OpenStudy (anonymous):

What are the approximate solutions of 4x2 + 2x = 17 rounded to the nearest hundredth? I dont understand this one.Please help

OpenStudy (jkristia):

You have quadratic equation there \[4x^2 + 2x - 17 = 0\] Then solve it using the quadratic formular \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] \[x = \frac{-2 \pm \sqrt{2^2-4(4)(-17)}}{2(4)}\]